OpenStudy (anonymous):

Find dy/dx when y=x^(4lnx)

OpenStudy (anonymous):

\[\large y=x^{4\ln x}=e^{\ln\left(x^{4\ln x}\right)}=e^{(4\ln x)\ln x}=e^{4(\ln x)^2}\] Use the chain rule: \[\large\begin{align*}\frac{dy}{dx}&=\frac{d}{dx}\left[e^{4(\ln x)^2}\right]\\\\ &=e^{4(\ln x)^2}\frac{d}{dx}\left[4(\ln x)^2\right]\\\\ &=x^{4\ln x}\frac{d}{dx}\left[4(\ln x)^2\right]\\\\ &=\cdots \end{align*}\]